OpenStudy (anonymous):

Help understanding a change in limits of integration: Original limits are from sqrt2 to 2. If u=sec(theta), then the new limits become pi/4 to pi/3, respectively. How did these numbers come about?

OpenStudy (anonymous):

I actually see where the pi/3 came from. Still having trouble seeing where pi/4 came from though...

OpenStudy (asnaseer):

\[u=\sec(\theta)=\frac{1}{\cos(\theta)}\]\[\therefore \theta=\cos^{-1}(\frac{1}{u})\]Now put in \(u=\sqrt{2}\) and \(u=2\) to get the new limits

OpenStudy (anonymous):

I was forgetting to rationalize the denominator during u=sqrt2 :3 Thanks for your help!

OpenStudy (asnaseer):

yw :)